OpenStudy (anonymous):

\[\sum_{k=2}^{3} (k ^{2}-2k)\] so confused

OpenStudy (amistre64):

what is the confusion about?

OpenStudy (amistre64):

the value at k=2, added to the value at k=3 .... seems simple enough

OpenStudy (anonymous):

I can't seem to find the sum of the series

OpenStudy (anonymous):

I have done it twice and still not getting the right answer

OpenStudy (amistre64):

show me your work then

OpenStudy (amistre64):

there are a few ways to approach this ... so knowing what way you tried will help determine your errors

OpenStudy (anonymous):

2^2-2(2)= 0 3^2-2(3)= 3 4^2-2(4)= 8 the sum would be 11 if 8+3 is added we get 11

OpenStudy (amistre64):

well, your adding is good, but im curious why you want to k=4? your index says k = 2 to 3

OpenStudy (anonymous):

oh crap I went to far didn't I

OpenStudy (amistre64):

over and beyond :)

OpenStudy (anonymous):

my bad

OpenStudy (anonymous):

so the answer would 3

OpenStudy (amistre64):

yes, with what you have posted i would say 3 is correct

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

youre welcome, good luck :)